### v1

# Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models

Abstract
The literature review is a crucial form of academic writing that involves complex processes of literature collection, organization, and summarization. The emergence of large language models (LLMs) has introduced promising tools to automate these processes. However, their actual capabilities in writing comprehensive literature reviews remain underexplored, such as whether they can generate accurate and reliable references. To address this gap, we propose a framework to assess the literature review writing ability of LLMs automatically. We evaluate the performance of LLMs across three tasks: generating references, writing abstracts, and writing literature reviews. We employ external tools for a multidimensional evaluation, which includes assessing hallucination rates in references, semantic coverage, and factual consistency with human-written context. By analyzing the experimental results, we find that, despite advancements, even the most sophisticated models still cannot avoid generating hallucinated references. Additionally, different models exhibit varying performance in literature review writing across different disciplines.
1 Introduction
The literature review is a form of academic writing designed to summarize, analyze, and evaluate existing research findings, theoretical perspectives, and methodologies. However, writing a literature review often requires extensive reading and summarizing of relevant literature. Literature reviews in popular fields may require tens or even hundreds of references, which is a complex and labor-intensive task. In recent years, deep learning models have been developed to assist in the automated writing of literature reviews Kontonatsios et al. (2020); Aliyu et al. (2018). The advent of large language models (LLMs) has brought promising advances in automating certain stages of literature review writing. With their powerful generative capabilities, enhanced by training on vast amounts of academic text, LLMs have emerged as a formidable tool for writing literature reviews Wang et al. (2024).
However, how well LLMs perform in writing literature reviews remains unclear. To the best of our knowledge, no comprehensive studies have yet evaluated the literature review writing capabilities of LLMs. Literature review writing requires, on one hand, deep familiarity with the studies in a specific field, and on the other hand, the ability to summarize the core contributions of that literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance LLMs’ domain-specific knowledge, most researchers use LLMs without RAG such as ChatGPT for literature review writing. Therefore, it is necessary to evaluate the literature review writing capabilities of these native LLMs.
Therefore, in this paper, we propose a framework for automatically assessing literature review writing ability in LLMs. The evaluation of LLMs’ ability to write literature reviews remains an open question. In this study, we use human-written literature reviews as the gold standard and employ external tools for a comprehensive evaluation. Specifically, we first collect human-written literature reviews to create the dataset. Then, we ask LLMs to complete three tasks: generating references, writing an abstract, and writing a literature review based on a given topic. The generated results are then evaluated from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
Our contribution can be summarized as follows.
  • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement. This framework encompasses multiple stages, including dataset construction, LLMs gnenrated data collection, and evaluation.
  • Second, we collect 1,106 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
  • Then, we evaluate the writing results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
  • Finally, we assess four LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs. Additionally, the performance of LLMs in writing literature reviews varies across different academic disciplines.
2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024) proposed a literature survey writing process for LLMs, called AutoSurvey. By RAG retrieving the latest relevant paper incorporates real-time knowledge. In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.
Additionally, more attention has been paid to the hallucination evaluation of reference citations. According to whether the output contradicts the prompt or not,  OpenAI et al. (2023) categorized hallucination phenomenon as open domain and closed-domain hallucinations. Closed-domain hallucination assessments require only comparisons with input contexts, whereas open-domain hallucination assessments require more extensive research Bubeck et al. (2023), such as human evaluation Ji et al. (2023).
Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generated 11 systematic reviews under the shoulder rotator cuff pathology topic, then to evaluate the hallucination rates of reference. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. For example,  Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rate. They first asked LLMs to generate 5 reference titles, then use BING Search API to label the reference automatically as true or not. The LLMs were then examined to see if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on the given 50 topics. ChatGPT generated 178 references in these proposals and then they employed five researchers to manually check these references and DOIs on Scopus, Google and PubMed search engines and found that among the references 109 references had valid DOIs. score  Aljamaan et al. (2024) given 5 medical topics, then used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generated 10 relevant references for each topic, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication data, link to the paper, and relevance of the reference to the keyword prompt. Then, they proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. The existing studies on evaluating citation hallucinations in literature reviews are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
3 Methodology
In this section, we present the framework for evaluating LLMs’ literature survey writing ability. The framework as shown in Figure 1, consists of three main stages: first, dataset construction; second, task design for evaluation; and third, assessment of the generated results.
3.1 Dataset Construction
As illustrated in Figure 1, we first collect literature survey data from the Annual Reviews website ( https://www.annualreviews.org/). Annual Reviews is an independent, nonprofit scholarly publishing company, which publishes 51 journals of review articles. The articles on the website are peer-reviewed, so they are of high quality and can be used as experimental materials. We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references. Then clean them as the experimental dataset. Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex. In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent.
Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$. For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
3.2 Task Design
We design three tasks as follows.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the 10 most relevant studies to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract according to the research topic provided in the title and the keywords. The length of the generated abstract is required to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies (with a total of 10 citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
Three task prompts are shown in Appendix Table LABEL:at1.
3.3 Evaluation Metrics
Based on the type of generated text, we divide the evaluation of the model’s results into two parts: first, the hallucination rate of the references generated by LLMs, and second, a comparison of the generated context with human-written results, including two dimensions: factual consistency and semantic coverage.
Figure 1: Illustration of the proposed framework.
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate accurate references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM.
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task1 and Task3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $C_{t}$ and $C_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $C_{t}$ and $C_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $c_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $c_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. therefore, when the other three elements in $r_{n}^{g}$ are correct, we also consider it as a reliable article.
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3. Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4. Finally, we defined the reference accuracy metric $Acc$ as Eq 4. A high accuracy metric indicates a low hallucination rate.
$Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
$Acc_{ref.}(j_{k})=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
$Acc.=\frac{1}{K}\sum_{k=0}^{K}{Acc_{ref.}(j_{k})}$ (4)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of each journal and each LLM, formulated as follows.
$S_{t}=1/(NMK)\sum_{K}\sum_{M}\sum_{N}s_{p_{i}}$ (5)
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from Semantic Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (6)
where $s$ is the number of correct generated reference titles for $p_{i}$. We search the generated title from an external academic search engine to estimate whether the title can be matched.
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{i}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
$Entail_{i}=\theta_{\text{NLI}}(a_{i},a_{i}^{g})=\begin{cases}1&\text{if }a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (7)
where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 89.
$Entail_{j_{k}}=\frac{1}{M}\sum_{i=0}^{M}{Entail_{i}}$ (8)
$Entail=\frac{1}{K}\sum_{k=0}^{K}{Entail_{j_{k}}}$ (9)
Additionally, we use commonly employed semantic similarity metrics to calculate the semantic coverage between the content generated by LLMs and human-written content. Specifically, we apply cosine similarity and the ROUGE metric for evaluation.
4 Experiments
4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1106 articles, the dataset distribution across journals as shown in Figure 7.
LLMs Selection. We select four LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, and Llama-3.2-3B-Instruct.
For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context. To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings.
4.2 Main Results
We present the results for the three tasks in Table 12, and 3. The performance of different models on each task is analyzed as follows.
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy score and title search rate, whereas Llama-3.2 records the lowest performance on both metrics for Task 1. Additionally, when evaluating the accuracy of LLM-generated references, treating the author item as correct if the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding highlights that accurately generating complete author lists remains a significant challenge for LLMs.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5 consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5 and human-written abstracts is 81.17%. Furthermore, Claude-3.5 demonstrates the highest factual consistency, with a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.77%. Additionally, we observe that Llama-3.2 achieves the highest score on the ROUGE-L metric, but does not show a clear advantage on other metrics. Given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text. In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
  Models Task 1
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 19.81 12.36 17.64
Llama-3.2-3B 14.36 3.39 6.86
GPT-4o 30.51 21.36 24.77
Claude-3.5-Sonnet 63.99 51.44 55.40
Table 1: The experimental results of the four LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 5, while “Acc.” represents the accuracy defined in Eq 4. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task 2
[…]
Qwen-2.5-72B 80.22 68.69 95.02 40.61 8.78 20.12
Llama-3.2-3B 79.28 61.90 92.14 40.35 8.96 20.52
GPT-4o 80.96 77.74 96.50 40.70 8.56 19.86
Claude-3.5-Sonnet 81.17 78.10 96.77 41.13 8.99 20.00
Table 2: Compare the performance of four LLMs on Task 2.
  Models Task 3
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 38.39 28.89 33.64
Llama-3.2-3B 21.00 4.85 7.27
GPT-4o 60.24 50.69 54.24
Claude-3.5-Sonnet 66.51 58.89 62.90
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 5.“Acc.” represents the accuracy defined in Eq 5. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task1 Task3
Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
Qwen-2.5-72B 7.18 12.57 13.24 12.13 17.31 17.88
Llama-3.2-3B 5.97 8.48 8.67 8.02 8.17 8.33
GPT-4o 8.66 18.78 19.50 18.33 28.81 29.63
Claude-3.5-Sonnet 16.05 24.34 25.21 21.15 33.25 34.41
Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3. “Title Acc.” refers to the title overlap between LLM-generated and human-written references, while “overlap” represents the article overlap between LLM-generated and human-written references. “first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
Figure 3: Task 1: Accuracy scores $\uparrow$ across different disciplines.
Figure 4: Task 2: NLI scores (TRUE)$\uparrow$ across different disciplines.
Figure 5: Task 3: Accuracy scores $\uparrow$ across different disciplines.
4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4o, and Qwen-2.5 shows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. The accuracy for journal name, page, and author is also relatively high. In contrast, Llama-3.2 shows higher accuracy in the page and author dimensions compared to other dimensions, but overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, GPT-4o, and Qwen-2.5. Notably, the accuracy of GPT-4o in certain dimensions approaches that of Claude-3.5. However, the performance of LLaMA-3.2 remains suboptimal.
4.4 Cross-Disciplinary Analysis
In this section, we compare the performance of LLMs across different disciplines. First, based on Dewey’s Decimal Classification, we categorize 51 journals into five disciplines: Biology, Chemistry, Mathematics, Physics, Social Science, and Technology. After categorization, there are 460 articles in the Biology category, 90 in Chemistry, 50 in Mathematics, 113 in Physics, 299 in Social Science, and 94 in Technology.
We then present bar charts in Figures  345, which illustrate the performance of different models across various tasks and disciplines.
First, we observe that in Task 1, as shown in Figure 3, almost all models exhibit the highest accuracy in the Mathematics discipline. Additionally, Claude-3.5 demonstrates relatively high accuracy in both Biology and Social science. GPT-4o, besides having the highest accuracy in Mathematics, shows higher accuracy in Physics and Social science compared to Biology and Technology. Qwen-2.5 also exhibits a relatively high accuracy in Physics, but it has the lowest accuracy in Chemistry. Llama-3.2 has the highest accuracy in Mathematics, but its accuracy in Technology is the lowest. Finally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5(f=7.65, p<.001), GPT-4o(f=11.87, p<.001), Qwen-2.5(f=12.39, p<.001), Llama-3.2(f=10.78, p<.001). These results indicate significant differences in performance across disciplines for each model.
Secondly, as shown in Figure 4, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. In terms of this metric, GPT-4o outperforms Claude-3.5 in Mathematics and Technology, indicating that while GPT-4o has a lower accuracy in reference generation, the difference between the two models in the abstract generation task is relatively small. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Llama-3.2 (f= 11.06, p<.001) Qwen2.5 (f= 16.88, p<.001) GPT-4o(f= 18.88, p<.001) Claude-3.5(f= 23.72, p<.001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Finally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure 4. It is evident that the accuracy of Claude-3.5 and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(f=33.33, p<.001), GPT-4o(f=39.09, p<.001), Qwen-2.5(f=32.84, p<.001), Llama-3.2(f=39.02, p<.001), indicating significant differences in each models’ performance across the various disciplines.
4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. Next, we provide a more detailed comparison of the references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
5 Conclusion
In this paper, we present a framework to assess the literature review writing abilities of LLMs. This framework includes three tasks designed to evaluate LLMs’ literature review writing capabilities. The generated outputs are then evaluated from multiple dimensions using various tools, such as Semantic Scholar and NLI models, focusing on aspects like hallucination rate, semantic coverage, and factual consistency compared to human-written texts. Finally, we analyze the performance of LLMs in writing literature reviews from the perspective of different academic disciplines.
We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references. This advantage is likely influenced by the training data of each model. Additionally, we observed that each model has different strengths across disciplines. Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology. However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references. This suggests that LLMs cite references during the writing process, which improves the authenticity of the references. Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.
6 Limitations
In this paper, we evaluate the ability of LLMs to write literature reviews. However, instead of assessing from conventional perspectives, such as the fluency of the generated context or topic coverage, we compare the model-generated results with human-written ones. Additionally, when processing the LLM-generated results, we often need to handle abbreviations of author names and journal titles. While we have made efforts to address these details thoroughly, some discrepancies may still exist.
Appendix A Prompts for Tasks
Table 5: Prompts for tasks.
Prompt Content
Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to find the 10 studies that are most relevant to the research topic provided in the "Title" and the "Keywords" below.
Please cite the studies according to the following JSON format. There is no need to provide any explanation before or after the JSON output. Ensure that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.". { "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
Title: title
Keywords: keywords
Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write an abstract according to the research topic provided in the "Title" and the "Keywords" below. Please write the abstract for about xx words, according to the JSON format as follows. There is no need to provide any explanation before or after the JSON output. {"Abstract": ""}
Title: title
Keywords: keywords
Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write a literature review according to the research topic provided in the "Title", “Abstract” and "Keywords" below. The literature review should be about 1000 words long. I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review). The output should be in JSON format as follows: { "Literature Review": "xxx", "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
The "Literature Review" field should be about 1000 words. The "References" field is a list of 10 references, and ensures that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
Title: title
Keywords: keywords
Abstract: abstract
Appendix B Data Distribution
Figure 7: Statistics of dataset.

### v2

@@
 This framework encompasses multiple stages, including dataset construction, LLMs gnenrated data collection, and evaluation.
-  • Second, we collect 1,106 literature reviews from 51 journals across 5 disciplines as the ground truth.
+  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth.
 We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
-  • Then, we evaluate the writing results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
+  • Then, we evaluate the generated results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
   • Finally, we assess four LLMs using the proposed framework.
 By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs.
-Additionally, the performance of LLMs in writing literature reviews varies across different academic disciplines.
+Additionally, the performance of LLMs in writing literature reviews varies across different disciplines.
 2 Related Work
@@
 In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.
+Wang_Hu_Wang_Yan_Sheng_He_2024 compared the use of LLMs across different academic tasks and found that researchers are more likely to use LLMs for literature review (LR) tasks than for paper understanding (PU) tasks.
 Additionally, more attention has been paid to the hallucination evaluation of reference citations.
@@
 Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$.
-For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
+For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles in $j_{k}$, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
 3.2 Task Design
 We design three tasks as follows.
-  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the 10 most relevant studies to the research topic.
+  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies to the research topic.
 Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page.
@@
 The literature survey should be about 1000 words long.
-LLMs also need to back up claims by citing previous studies (with a total of 10 citations in the literature survey).
+LLMs also need to back up claims by citing previous studies (with a total of $N$ citations in the literature survey).
 In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@
 Finally, we defined the reference accuracy metric $Acc$ as Eq 4.
-A high accuracy metric indicates a low hallucination rate.
+A higher accuracy metric indicates a lower hallucination rate.
 $Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
@@
 where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 8.
-$Entail_{j_{k}}=\frac{1}{M}\sum_{i=0}^{M}{Entail_{i}}$ (8)
-$Entail=\frac{1}{K}\sum_{k=0}^{K}{Entail_{j_{k}}}$ (9)
-Additionally, we use commonly employed semantic similarity metrics to calculate the semantic coverage between the content generated by LLMs and human-written content.
-Specifically, we apply cosine similarity and the ROUGE metric for evaluation.
+$Entail=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}{Entail_{i}}$ (8)
+Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context.
+Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation.
+For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
+KPR, first proposed by  Qi_Xu_Guo_Wang_Zhang_Xu_2024, is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents.
+Since the human-written literature review is lengthy and difficult to compare directly, we adapt the KPR metric by first using GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculating the coverage of these key points by the model-generated literature review as Eq.
+9.
+$KPR=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (9)
+where $c_{i}^{g}$ denotes the literature review generated by LLMs. Finally, we also concatenate key points and compute the ROUGE metric between the key point and $c_{i}^{g}$.
 4 Experiments
 4.1 Experimental Settings
-Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1106 articles, the dataset distribution across journals as shown in Figure 7.
-LLMs Selection. We select four LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, and Llama-3.2-3B-Instruct.
+Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
+LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3 and Llama-3.2-3B-Instruct.
+In task 1 and task 3, we set $N$ as 10, each model generates 10 references.
 For the generated reference evaluation, we use Semantic Scholar as the external database.
@@
 (2024).
-So, we employ TRUE Honovich et al.
-(2022), along with GPT-4o as NLI models for evaluating factual consistency in context.
-To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings.
+So, for Task 2, we employ TRUE Honovich et al.
+(2022), along with GPT-4o as NLI models for evaluating factual consistency in context; To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings.
+For task 3, we employ GPT-4o as the NLI model in Eq 9, and set $q$ as 10.
 4.2 Main Results
@@
 Results for Task 1.
-By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy score and title search rate, whereas Llama-3.2 records the lowest performance on both metrics for Task 1.
-Additionally, when evaluating the accuracy of LLM-generated references, treating the author item as correct if the first author is identified correctly results in a 3–5% increase in accuracy scores across all models.
-This finding highlights that accurately generating complete author lists remains a significant challenge for LLMs.
+By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy and title search rate, whereas Llama-3.2 exhibits the lowest performance on both metrics for Task 1.
+Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models.
+This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs.
+Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2.
+The graph shows that the majority of the correctly cited references are concentrated between 2010 and 2020.
 Results for Task 2.
@@
 This suggests that the generated references and the literature review text impose constraints on each other.
+On the other hand, we find that Claude-3.5 achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5 recalls the most claims from the human-written literature review.
+While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
   Models Task 1
 $\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
-Qwen-2.5-72B 19.81 12.36 17.64
-Llama-3.2-3B 14.36 3.39 6.86
-GPT-4o 30.51 21.36 24.77
-Claude-3.5-Sonnet 63.99 51.44 55.40
+Qwen-2.5-72B 21.80 12.25 17.58
+Llama-3.2-3B 16.62 3.45 6.95
+DeepSeek-V3 56.04 46.33 50.66
+GPT-4o 32.07 21.65 24.65
+Claude-3.5-Sonnet 64.82 51.59 55.77
 Table 1: The experimental results of the four LLMs in Task 1.
@@
 […]
-Qwen-2.5-72B 80.22 68.69 95.02 40.61 8.78 20.12
-Llama-3.2-3B 79.28 61.90 92.14 40.35 8.96 20.52
-GPT-4o 80.96 77.74 96.50 40.70 8.56 19.86
-Claude-3.5-Sonnet 81.17 78.10 96.77 41.13 8.99 20.00
+Qwen-2.5-72B 80.22 69.52 95.02 40.61 8.78 20.12
+Llama-3.2-3B 79.28 62.39 92.14 40.35 8.96 20.52
+DeepSeek-V3 80.96 78.55 96.84 41.13 8.98 20.33
+GPT-4o 80.96 77.91 96.50 40.70 8.56 19.86
+Claude-3.5-Sonnet 81.17 78.90 96.77 41.13 8.99 20.00
 Table 2: Compare the performance of four LLMs on Task 2.
   Models Task 3
-$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
-Qwen-2.5-72B 38.39 28.89 33.64
-Llama-3.2-3B 21.00 4.85 7.27
-GPT-4o 60.24 50.69 54.24
-Claude-3.5-Sonnet 66.51 58.89 62.90
+References Literature Review
+[…]
+Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
+Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
+DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
+GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
+Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
 Table 3: The experimental results of the four LLMs in Task 3.
 “$S_{t}$” refers to the title retrieval rate as defined in Eq 5.“Acc.” represents the accuracy defined in Eq 5.
-“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
+“KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
   Models Task1 Task3
 Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
-Qwen-2.5-72B 7.18 12.57 13.24 12.13 17.31 17.88
-Llama-3.2-3B 5.97 8.48 8.67 8.02 8.17 8.33
-GPT-4o 8.66 18.78 19.50 18.33 28.81 29.63
-Claude-3.5-Sonnet 16.05 24.34 25.21 21.15 33.25 34.41
+Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
+Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
+DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
+GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
+Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
 Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3.
@@
 “first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
-Figure 3: Task 1: Accuracy scores $\uparrow$ across different disciplines.
-Figure 4: Task 2: NLI scores (TRUE)$\uparrow$ across different disciplines.
-Figure 5: Task 3: Accuracy scores $\uparrow$ across different disciplines.
+Figure 2: Distribution of LLM-generated true references over years.
 4.3 Analyze LLM-Generated References from Different Dimensions
@@
 As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1.
-Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4o, and Qwen-2.5 shows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
-The accuracy for journal name, page, and author is also relatively high.
-In contrast, Llama-3.2 shows higher accuracy in the page and author dimensions compared to other dimensions, but overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
-Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, GPT-4o, and Qwen-2.5.
-Notably, the accuracy of GPT-4o in certain dimensions approaches that of Claude-3.5.
+Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4, and Qwen-2.5 follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
+Accuracy for journal name, page, and author is also relatively high.
+However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions.
+In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions.
+However, overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
+Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, DeepSeek-V3, GPT-4o, and Qwen-2.5, with particularly obvious gains in the author dimension.
+The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
+Notably, the accuracy of DeepSeek-V3 and GPT-4 in certain dimensions approaches or even exceeds that of Claude-3.5.
 However, the performance of LLaMA-3.2 remains suboptimal.
@@
 After categorization, there are 460 articles in the Biology category, 90 in Chemistry, 50 in Mathematics, 113 in Physics, 299 in Social Science, and 94 in Technology.
-We then present bar charts in Figures  345, which illustrate the performance of different models across various tasks and disciplines.
-First, we observe that in Task 1, as shown in Figure 3, almost all models exhibit the highest accuracy in the Mathematics discipline.
-Additionally, Claude-3.5 demonstrates relatively high accuracy in both Biology and Social science.
-GPT-4o, besides having the highest accuracy in Mathematics, shows higher accuracy in Physics and Social science compared to Biology and Technology.
-Qwen-2.5 also exhibits a relatively high accuracy in Physics, but it has the lowest accuracy in Chemistry.
-Llama-3.2 has the highest accuracy in Mathematics, but its accuracy in Technology is the lowest.
-Finally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences.
-The results are as follows: Claude-3.5(f=7.65, p<.001), GPT-4o(f=11.87, p<.001), Qwen-2.5(f=12.39, p<.001), Llama-3.2(f=10.78, p<.001).
-These results indicate significant differences in performance across disciplines for each model.
-Secondly, as shown in Figure 4, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2.
+We then present bar charts in Figures  LABEL:f4, which illustrate the performance of different models across various tasks and disciplines.
+First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline.
+Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences.
+The results are as follows: Claude-3.5 (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5 (p<.0001), Llama-3.2 (p=0.065).
+These results indicate significant differences in performance across disciplines for each model except Llama-3.2.
+Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2.
 We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology.
-In terms of this metric, GPT-4o outperforms Claude-3.5 in Mathematics and Technology, indicating that while GPT-4o has a lower accuracy in reference generation, the difference between the two models in the abstract generation task is relatively small.
 Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines.
-The results are as follows: Llama-3.2 (f= 11.06, p<.001) Qwen2.5 (f= 16.88, p<.001) GPT-4o(f= 18.88, p<.001) Claude-3.5(f= 23.72, p<.001).
+The results are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5 (p<.001), Llama-3.2 (p<.0001).
 These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
-Finally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure 4.
-It is evident that the accuracy of Claude-3.5 and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines.
+Additionally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c.
+It is evident that the accuracy of Claude-3.5, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines.
 Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology.
-The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(f=33.33, p<.001), GPT-4o(f=39.09, p<.001), Qwen-2.5(f=32.84, p<.001), Llama-3.2(f=39.02, p<.001), indicating significant differences in each models’ performance across the various disciplines.
+The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5(p<.0001), Llama-3.2(p<.001), indicating significant differences in each models’ performance across the various disciplines.
+Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d.
+The results from the figure indicate that the differences between models—Claude-3.5, DeepSeek-V3, GPT-4, and Qwen-2.5—are not significant across various disciplines, a finding that is also supported by statistical tests.
+Specifically, the p-values are as follows: Claude-3.5 (p=0.46), DeepSeek-V3 (p=0.23), GPT-4 (p=0.18), Qwen-2.5 (p=0.10), and Llama-3.2 (p<0.001).
 4.5 Comparison of LLM-Cited and Human-cited References
@@
 For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1.
-The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
+This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
 5 Conclusion
@@
 Finally, we analyze the performance of LLMs in writing literature reviews from the perspective of different academic disciplines.
-We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references.
+We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references.
 This advantage is likely influenced by the training data of each model.

### v3

@@
-Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models
+Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition
 Abstract
-The literature review is a crucial form of academic writing that involves complex processes of literature collection, organization, and summarization.
-The emergence of large language models (LLMs) has introduced promising tools to automate these processes.
-However, their actual capabilities in writing comprehensive literature reviews remain underexplored, such as whether they can generate accurate and reliable references.
-To address this gap, we propose a framework to assess the literature review writing ability of LLMs automatically.
-We evaluate the performance of LLMs across three tasks: generating references, writing abstracts, and writing literature reviews.
-We employ external tools for a multidimensional evaluation, which includes assessing hallucination rates in references, semantic coverage, and factual consistency with human-written context.
-By analyzing the experimental results, we find that, despite advancements, even the most sophisticated models still cannot avoid generating hallucinated references.
-Additionally, different models exhibit varying performance in literature review writing across different disciplines.
+Large language models (LLMs) have emerged as a potential solution to automate the complex processes involved in writing literature reviews, such as literature collection, organization, and summarization.
+However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews.
+This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature writing: reference generation, literature summary, and literature review composition.
+We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts.
+The experimental results reveal that even the most advanced models still generate hallucinated references, despite recent progress.
+Moreover, we observe that the performance of different models varies across disciplines when it comes to writing literature reviews.
+These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews.
 1 Introduction
-The literature review is a form of academic writing designed to summarize, analyze, and evaluate existing research findings, theoretical perspectives, and methodologies.
-However, writing a literature review often requires extensive reading and summarizing of relevant literature.
-Literature reviews in popular fields may require tens or even hundreds of references, which is a complex and labor-intensive task.
-In recent years, deep learning models have been developed to assist in the automated writing of literature reviews Kontonatsios et al.
+The literature review is a critical component of academic writing that aims to synthesize, critique, and assess the current state of knowledge in a particular field.
+It involves a comprehensive examination of published research articles, theoretical frameworks, and research methodologies related to a specific topic.
+Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds.
+To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews.
+Kontonatsios et al.
 (2020); Aliyu et al.
 (2018).
-The advent of large language models (LLMs) has brought promising advances in automating certain stages of literature review writing.
-With their powerful generative capabilities, enhanced by training on vast amounts of academic text, LLMs have emerged as a formidable tool for writing literature reviews Wang et al.
+The emergence of large language models (LLMs) has introduced a promising avenue for automating various aspects of the literature review writing process.
+LLMs, trained on extensive collections of academic texts, have developed impressive generative capabilities that make them a potent tool for assisting researchers in crafting literature reviews.
+By leveraging the knowledge acquired from vast amounts of scholarly literature, these models can help researchers quickly identify relevant sources, summarize key findings, and generate coherent syntheses of the existing body of work Wang et al.
+(2024b); Agarwal et al.
 (2024).
-However, how well LLMs perform in writing literature reviews remains unclear.
-To the best of our knowledge, no comprehensive studies have yet evaluated the literature review writing capabilities of LLMs.
-Literature review writing requires, on one hand, deep familiarity with the studies in a specific field, and on the other hand, the ability to summarize the core contributions of that literature.
-While techniques such as Retrieval-Augmented Generation (RAG) can enhance LLMs’ domain-specific knowledge, most researchers use LLMs without RAG such as ChatGPT for literature review writing.
-Therefore, it is necessary to evaluate the literature review writing capabilities of these native LLMs.
-Therefore, in this paper, we propose a framework for automatically assessing literature review writing ability in LLMs.
-The evaluation of LLMs’ ability to write literature reviews remains an open question.
-In this study, we use human-written literature reviews as the gold standard and employ external tools for a comprehensive evaluation.
-Specifically, we first collect human-written literature reviews to create the dataset.
-Then, we ask LLMs to complete three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
-The generated results are then evaluated from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
+The process of writing a literature review requires a deep understanding of the studies within a particular field, as well as the ability to summarize the core contributions of the relevant literature.
+While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al.
+(2024a).
+Consequently, it is crucial to evaluate the performance of these native LLMs in the context of literature review writing to determine their effectiveness and limitations.
+Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation.
+We first collect a dataset of human-written literature reviews to serve as a benchmark for evaluating the performance of LLMs.
+We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic.
+Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
+By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
 Our contribution can be summarized as follows.
   • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement.
-This framework encompasses multiple stages, including dataset construction, LLMs gnenrated data collection, and evaluation.
+This framework encompasses multiple stages, including the compilation of a literature review dataset construction, the collection of LLM-generated output, and the evaluation of LLM performance.
   • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth.
-We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
-  • Then, we evaluate the generated results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
-  • Finally, we assess four LLMs using the proposed framework.
-By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs.
-Additionally, the performance of LLMs in writing literature reviews varies across different disciplines.
+We then design three tasks for accessing LLMs in literature writing: reference generation, literature summary, and literature composition on a given topic.
+  • Then, we evaluate the generated results of LLMs from multiple perspectives, including the hallucination rate in generated references, factual consistency, and semantic coverage compared to human-written content.
+  • Finally, we assess five LLMs using the proposed framework.
+By analyzing the experimental results, we find that hallucinated references remain a prevalent issue for current LLMs.
+Furthermore, the performance of LLMs in writing literature reviews varies across different disciplines.
 2 Related Work
@@
 For example,  Wang et al.
-(2024) proposed a literature survey writing process for LLMs, called AutoSurvey.
-By RAG retrieving the latest relevant paper incorporates real-time knowledge.
-In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.
-Wang_Hu_Wang_Yan_Sheng_He_2024 compared the use of LLMs across different academic tasks and found that researchers are more likely to use LLMs for literature review (LR) tasks than for paper understanding (PU) tasks.
-Additionally, more attention has been paid to the hallucination evaluation of reference citations.
-According to whether the output contradicts the prompt or not,  OpenAI et al.
-(2023) categorized hallucination phenomenon as open domain and closed-domain hallucinations.
-Closed-domain hallucination assessments require only comparisons with input contexts, whereas open-domain hallucination assessments require more extensive research Bubeck et al.
-(2023), such as human evaluation Ji et al.
-(2023).
-Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generated 11 systematic reviews under the shoulder rotator cuff pathology topic, then to evaluate the hallucination rates of reference.
+(2024b) proposed a literature survey writing process for LLMs, called AutoSurvey, which incorporates real-time knowledge by retrieving the latest relevant papers using a retrieval-augmented generation (RAG) approach.
+Similarly,  Agarwal et al.
+(2024) investigated the zero-shot capabilities of recent LLMs in generating literature reviews from abstracts.
+They employed a two-step approach: first retrieving relevant papers via an external knowledge base, then outlining and executing a structured plan to generate the review.
+The authors focused on the LLMs’ ability to effectively summarize the literature.
+Additionally, for evaluating the literature review writing capabilities of LLMs, more attention has been paid to assessing hallucinations in reference citations.
+For example, Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generate 11 systematic reviews on shoulder rotator cuff pathology and evaluated the hallucination rates of the references.
 They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4.
-For example,  Agrawal et al.
-(2024) chose 200 computer science topics to evaluate LLMs’ hallucination rate.
-They first asked LLMs to generate 5 reference titles, then use BING Search API to label the reference automatically as true or not.
-The LLMs were then examined to see if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
+Similarly, Agrawal et al.
+(2024) chose 200 computer science topics to evaluate LLMs’ hallucination rates.
+They first asked LLMs to generate 5 reference titles, then used the BING Search API to automatically label the references as true or not.
+The LLMs were then examined to determine if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
 Athaluri et al.
 (2023) assessed the hallucinations of LLMs in research proposal writing.
-They used ChatGPT to generate 50 research proposals based on the given 50 topics.
-ChatGPT generated 178 references in these proposals and then they employed five researchers to manually check these references and DOIs on Scopus, Google and PubMed search engines and found that among the references 109 references had valid DOIs.
-score  Aljamaan et al.
-(2024) given 5 medical topics, then used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generated 10 relevant references for each topic, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication data, link to the paper, and relevance of the reference to the keyword prompt.
-Then, they proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date.
-The existing studies on evaluating citation hallucinations in literature reviews are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance.
+They used ChatGPT to generate 50 research proposals based on 50 given topics.
+ChatGPT generated 178 references in these proposals, and then they employed five researchers to manually check these references and DOIs on Scopus, Google, and PubMed search engines.
+They found that among the references, 109 had valid DOIs.
+Moreover, Aljamaan et al.
+(2024) provided 5 medical topics and used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication date, link to the paper, and relevance of the reference to the keyword prompt.
+They then proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date.
+While these studies are very useful attempts at evaluating citation hallucinations in literature reviews, they are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance.
 Moreover, these evaluations often involve manual human intervention.
 3 Methodology
-In this section, we present the framework for evaluating LLMs’ literature survey writing ability.
-The framework as shown in Figure 1, consists of three main stages: first, dataset construction; second, task design for evaluation; and third, assessment of the generated results.
+In this section, we propose a framework for evaluating LLMs’ literature survey writing ability.
+The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
 3.1 Dataset Construction
-As illustrated in Figure 1, we first collect literature survey data from the Annual Reviews website ( https://www.annualreviews.org/).
-Annual Reviews is an independent, nonprofit scholarly publishing company, which publishes 51 journals of review articles.
-The articles on the website are peer-reviewed, so they are of high quality and can be used as experimental materials.
-We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references.
-Then clean them as the experimental dataset.
 Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex.
 In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent.
-Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$.
-For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles in $j_{k}$, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
+As illustrated in Figure 1, we first collect publicly available information of literature reviews (i.e., the title, authors, abstract, and keywords) from the Annual Reviews website (https://www.annualreviews.org/).
+Annual Reviews, an independent nonprofit publisher, produces 51 review journals spanning various scientific disciplines.
+Invited experts write comprehensive, authoritative reviews that synthesize and summarize the most significant primary research literature in their field, providing a valuable resource for researchers to stay current with the latest developments.
+We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references, and then clean them to create the experimental dataset.
+Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles.
+Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
 3.2 Task Design
-We design three tasks as follows.
-  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies to the research topic.
+Since literature review writing primarily involves the collection and synthesis of relevant research, we design three tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
+  • Task 1: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic.
 Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page.
 In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
-  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract according to the research topic provided in the title and the keywords.
-The length of the generated abstract is required to be the same as the length of the original abstract of the article.
+  • Task 2: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract $a_{i}^{g}$according to the research topic provided in the title and the keywords.
+The length of the generated abstract is to be the same as the length of the original abstract of the article.
 In this task, we evaluate whether LLMs can write an abstract based on the given topic.
-  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract.
+  • Task 3: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract.
 The literature survey should be about 1000 words long.
-LLMs also need to back up claims by citing previous studies (with a total of $N$ citations in the literature survey).
+LLMs also need to back up claims by citing previous studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature survey).
 In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@
 Based on the type of generated text, we divide the evaluation of the model’s results into two parts: first, the hallucination rate of the references generated by LLMs, and second, a comparison of the generated context with human-written results, including two dimensions: factual consistency and semantic coverage.
-Figure 1: Illustration of the proposed framework.
+Figure 1: Illustration of the experiment process.
 Reference hallucination evaluation metrics.
-Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate accurate references.
+Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references.
 In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM.
-For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task1 and Task3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the article title, authors, journal, year, volume, first page, and last page.
+A higher accuracy metric indicates a lower hallucination rate.
+For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task 1 and Task 3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the generated article title, authors, journal, year, volume, first page, and last page.
 Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
@@
 First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines.
-Then, we obtain two candidate sets $C_{t}$ and $C_{a}$ respectively.
-Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $C_{t}$ and $C_{a}$.
+Then, we obtain two candidate sets $Z_{t}$ and $Z_{a}$ respectively.
+Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $Z_{t}$ and $Z_{a}$.
 For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$.
-Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $c_{n}^{g}$ of $r_{n}^{g}$.
-Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $c_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1).
+Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $z_{n}^{g}$ of $r_{n}^{g}$.
+Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $z_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1).
 On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article.
@@
 On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article.
-therefore, when the other three elements in $r_{n}^{g}$ are correct, we also consider it as a reliable article.
+Therefore, when the other three elements in $r_{n}^{g}$ are matching, we also consider it as a reliable article.
 $\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
 For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2.
-Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3.
-Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4.
-Finally, we defined the reference accuracy metric $Acc$ as Eq 4.
-A higher accuracy metric indicates a lower hallucination rate.
+Then, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(p_{i})$ across articles, as shown in Eq 3.
 $Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
-$Acc_{ref.}(j_{k})=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
-$Acc.=\frac{1}{K}\sum_{k=0}^{K}{Acc_{ref.}(j_{k})}$ (4)
+$Acc.=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
 On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness.
@@
 (2024)’s work, they determined by ground-truth labels assigned using the Bing search API.
-Encouraged by this work, we also calculate the title search scores of each journal and each LLM, formulated as follows.
-$S_{t}=1/(NMK)\sum_{K}\sum_{M}\sum_{N}s_{p_{i}}$ (5)
-$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from Semantic Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (6)
-where $s$ is the number of correct generated reference titles for $p_{i}$. We search the generated title from an external academic search engine to estimate whether the title can be matched.
+Encouraged by this work, we also calculate the title search scores of generated references for each LLM, formulated as follows.
+$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}$ (4)
+$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (5)
+where $s_{p_{i}}$ is the number of the generated title that can be matched in the external scholar base.
+We search the generated title from an external academic search engine to estimate whether the title can be matched.
 Context evaluation metrics.
@@
 Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text.
-For example, we calculate the NLI score $Entail_{i}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
-$Entail_{i}=\theta_{\text{NLI}}(a_{i},a_{i}^{g})=\begin{cases}1&\text{if}a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (7)
+For example, we calculate the NLI score $Entail_{p_{i}}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
+$Entail_{p_{i}}=\theta_{\text{NLI}}(a_{i}^{g},a_{i})=\begin{cases}1&\text{if }a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (6)
 where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 7.
-$Entail=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}{Entail_{i}}$ (8)
+$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}$ (7)
 Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context.
@@
 For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
-KPR, first proposed by  Qi_Xu_Guo_Wang_Zhang_Xu_2024, is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents.
-Since the human-written literature review is lengthy and difficult to compare directly, we adapt the KPR metric by first using GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculating the coverage of these key points by the model-generated literature review as Eq.
-9.
-$KPR=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (9)
-where $c_{i}^{g}$ denotes the literature review generated by LLMs. Finally, we also concatenate key points and compute the ROUGE metric between the key point and $c_{i}^{g}$.
+KPR, first proposed by  Qi et al.
+(2024), is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents.
+Since human-written literature reviews are lengthy and difficult to compare directly, we adopt the KPR method to measure the extent to which LLM-generated content covers the key points in human-written literature reviews.
+Specifically, we first use GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculate the coverage of these key points by the model-generated literature review as Eq.
+8.
+$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (8)
+where $c_{i}^{g}$ denotes the literature review generated by LLMs.
+Finally, we also concatenate key points and compute the ROUGE metric between the key points and $c_{i}^{g}$.
 4 Experiments
@@
 Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
-LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3 and Llama-3.2-3B-Instruct.
+LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
 In task 1 and task 3, we set $N$ as 10, each model generates 10 references.
@@
 Results for Task 1.
-By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy and title search rate, whereas Llama-3.2 exhibits the lowest performance on both metrics for Task 1.
+By analyzing the results in Table 1, we observe that Claude-3.5-Sonnet achieves the highest accuracy and title search rate, whereas Llama-3.2-3B exhibits the lowest performance on both metrics for Task 1.
 Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models.
@@
 Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2.
-The graph shows that the majority of the correctly cited references are concentrated between 2010 and 2020.
+The graph shows that the majority of the correctly cited references for almost all LLMs involved in the task are concentrated between 2010 and 2020.
 Results for Task 2.
 We use various metrics to evaluate model performance.
-As shown in Table 2, Claude-3.5 consistently achieves the best overall performance across nearly all metrics.
-The average semantic similarity between the abstracts generated by Claude-3.5 and human-written abstracts is 81.17%.
-Furthermore, Claude-3.5 demonstrates the highest factual consistency, with a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.77%.
-Additionally, we observe that Llama-3.2 achieves the highest score on the ROUGE-L metric, but does not show a clear advantage on other metrics.
-Given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
+As shown in Table 2, Claude-3.5-Sonnet consistently achieves the best overall performance across nearly all metrics.
+The average semantic similarity between the abstracts generated by Claude-3.5-Sonnet and human-written abstracts is 81.17%.
+Furthermore, Claude-3.5-Sonnet and DeepSeek-V3 exhibit the highest factual consistency, achieving a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.84%, respectively.
+Additionally, we observe that Llama-3.2-3B achieves the highest score on the ROUGE-L metric but does not exhibit a clear advantage on other metrics.
+This suggests that, given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
 Results for Task 3.
 As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3.
-Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text.
+Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text (Gao et al., 2023).
 In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly.
 This suggests that the generated references and the literature review text impose constraints on each other.
-On the other hand, we find that Claude-3.5 achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5 recalls the most claims from the human-written literature review.
+On the other hand, we find that Claude-3.5-Sonnet achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5-Sonnet recalls the most claims from the human-written literature review.
 While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
@@
 Claude-3.5-Sonnet 64.82 51.59 55.77
-Table 1: The experimental results of the four LLMs in Task 1.
+Table 1: The experimental results of the five LLMs in Task 1.
 “$S_{t}$” refers to the title search rate as defined in Eq 4, while “Acc.” represents the accuracy defined in Eq 3.
@@
 In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2.
-As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1.
-Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4, and Qwen-2.5 follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
+As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in Task 1.
+Additionally, the accuracy of reference generation in Task 1 for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
 Accuracy for journal name, page, and author is also relatively high.
@@
 In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions.
-However, overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
-Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, DeepSeek-V3, GPT-4o, and Qwen-2.5, with particularly obvious gains in the author dimension.
+However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
+Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension.
 The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
-Notably, the accuracy of DeepSeek-V3 and GPT-4 in certain dimensions approaches or even exceeds that of Claude-3.5.
-However, the performance of LLaMA-3.2 remains suboptimal.
+Notably, the accuracy of DeepSeek-V3 and GPT-4o in certain dimensions approaches or even exceeds that of Claude-3.5-Sonnet.
+However, the performance of LLaMA-3.2-3B remains suboptimal.
 4.4 Cross-Disciplinary Analysis
@@
 Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences.
-The results are as follows: Claude-3.5 (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5 (p<.0001), Llama-3.2 (p=0.065).
-These results indicate significant differences in performance across disciplines for each model except Llama-3.2.
+The results are as follows: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
+These results indicate significant differences in performance across disciplines for each model except Llama-3.2-3B.
 Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2.
@@
 Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines.
-The results are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5 (p<.001), Llama-3.2 (p<.0001).
+The results are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
 These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
-Additionally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c.
-It is evident that the accuracy of Claude-3.5, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines.
+Additionally, we examine the references accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c.
+It is evident that the accuracy of Claude-3.5-Sonnet, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines.
 Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology.
-The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5(p<.0001), Llama-3.2(p<.001), indicating significant differences in each models’ performance across the various disciplines.
+The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001), indicating significant differences in each models’ performance across the various disciplines.
 Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d.
-The results from the figure indicate that the differences between models—Claude-3.5, DeepSeek-V3, GPT-4, and Qwen-2.5—are not significant across various disciplines, a finding that is also supported by statistical tests.
-Specifically, the p-values are as follows: Claude-3.5 (p=0.46), DeepSeek-V3 (p=0.23), GPT-4 (p=0.18), Qwen-2.5 (p=0.10), and Llama-3.2 (p<0.001).
+The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests.
+Specifically, the p-values are as follows: Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
 4.5 Comparison of LLM-Cited and Human-cited References
@@
 Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article.
-As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate.
+As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5-Sonnet are also cited in the human-written article, indicating a relatively high overlap rate.
 Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3.
 Notably, GPT-4o models show a 10% increase in the overlap rate.
-Next, we provide a more detailed comparison of the references across various dimensions.
-As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower.
-For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1.
-This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
+We provide a more detailed comparison of the references across various dimensions in Appendix C.
+4.6 Human Evaluation
+To evaluate the reliability of our automatic assessment method for identifying hallucinated references, we conduct a comparative analysis involving 100 LLM-generated references.
+These references were assessed by three annotators and the final manual results were obtained by majority vote.
+The results demonstrated a kappa agreement of 0.71 between the automatic and human assessments, signifying a relatively high level of consistency and supporting the reliability of our method.
+Furthermore, when using human assessment results as the gold standard, the automatic assessment method achieved an accuracy of 86%, further validating its effectiveness.
 5 Conclusion
@@
 Finally, we analyze the performance of LLMs in writing literature reviews from the perspective of different academic disciplines.
-We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references.
-This advantage is likely influenced by the training data of each model.
-Additionally, we observed that each model has different strengths across disciplines.
-Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology.
-However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
-When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references.
-This suggests that LLMs cite references during the writing process, which improves the authenticity of the references.
-Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
-In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.
-6 Limitations
+Limitations
 In this paper, we evaluate the ability of LLMs to write literature reviews.
@@
 Abstract: abstract
+Figure 4: Statistics of dataset.
 Appendix B Data Distribution
-Figure 4: Statistics of dataset.
+Statistics of the dataset are shown in Figure 4.
+Appendix C Comparison of LLM-Cited and Human-cited References from Different Dimensions
+We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions.
+As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower.
+For Task 3, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1.
+This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
+Appendix D Discussion
+We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references.
+This advantage is likely influenced by the training data of each model.
+Additionally, we observed that each model has different strengths across disciplines.
+Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology.
+However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
+When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references.
+This suggests that LLMs cite references during the writing process, which improves the authenticity of the references.
+Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
+In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.

### v4

@@
 However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews.
-This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature writing: reference generation, literature summary, and literature review composition.
+This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature review writing: reference generation, abstract writing, and literature review composition.
 We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts.
@@
 These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews.
+The dataset and code used in this study are publicly available at an anonymous repository  [^1] https://anonymous.4open.science/r/Eval_LLM_LR-C657.
+Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition
+Xuemei Tang1 Xufeng Duan1 Zhenguang G. Cai1,2 1Department of Linguistics and Modern Languages, The Chinese University of Hong Kong 2Brain and Mind Institute, The Chinese University of Hong Kong
 1 Introduction
@@
 Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds.
-To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews.
-Kontonatsios et al.
-(2020); Aliyu et al.
+To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews Aliyu et al.
+(2018); Kontonatsios et al.
 (2020).
-The emergence of large language models (LLMs) has introduced a promising avenue for automating various aspects of the literature review writing process.
-LLMs, trained on extensive collections of academic texts, have developed impressive generative capabilities that make them a potent tool for assisting researchers in crafting literature reviews.
-By leveraging the knowledge acquired from vast amounts of scholarly literature, these models can help researchers quickly identify relevant sources, summarize key findings, and generate coherent syntheses of the existing body of work Wang et al.
+The emergence of large language models (LLMs) has introduced a promising avenue for automating key aspects of literature review writing, including identifying relevant sources, summarizing findings, and generating coherent syntheses Wang et al.
 (2024b); Agarwal et al.
+(2024); Hsu et al.
 (2024).
-The process of writing a literature review requires a deep understanding of the studies within a particular field, as well as the ability to summarize the core contributions of the relevant literature.
 While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al.
 (2024a).
-Consequently, it is crucial to evaluate the performance of these native LLMs in the context of literature review writing to determine their effectiveness and limitations.
+Consequently, it is crucial to evaluate the performance of these naive LLMs in the context of literature review writing to determine their effectiveness and limitations.
 Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation.
@@
 We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic.
-Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
+Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated abstract and literature review compared to the human-written context.
 By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
@@
   • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth.
-We then design three tasks for accessing LLMs in literature writing: reference generation, literature summary, and literature composition on a given topic.
+We then design three tasks for accessing LLMs in literature writing: reference generation, abstract writing, and literature composition on a given topic.
   • Then, we evaluate the generated results of LLMs from multiple perspectives, including the hallucination rate in generated references, factual consistency, and semantic coverage compared to human-written content.
@@
 2 Related Work
-Recent studies have increasingly focused on using LLMs for writing literature reviews.
+Recent studies have explored LLMs for literature review generation.
 For example, Wang et al.
-(2024b) proposed a literature survey writing process for LLMs, called AutoSurvey, which incorporates real-time knowledge by retrieving the latest relevant papers using a retrieval-augmented generation (RAG) approach.
-Similarly,  Agarwal et al.
-(2024) investigated the zero-shot capabilities of recent LLMs in generating literature reviews from abstracts.
-They employed a two-step approach: first retrieving relevant papers via an external knowledge base, then outlining and executing a structured plan to generate the review.
-The authors focused on the LLMs’ ability to effectively summarize the literature.
-Additionally, for evaluating the literature review writing capabilities of LLMs, more attention has been paid to assessing hallucinations in reference citations.
-For example, Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generate 11 systematic reviews on shoulder rotator cuff pathology and evaluated the hallucination rates of the references.
-They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4.
+(2024b) proposed AutoSurvey, which incorporates up-to-date papers via retrieval-augmented generation.
+Agarwal et al.
+(2024) examined zero-shot LLM review generation using a two-step retrieval and outlining process.
+More recently, Liang et al.
+(2025) presented SurveyX, an efficient system that optimizes retrieval, extraction, and outline generation, supporting multimodal outputs such as figures and tables.
+Additionally, recent efforts to evaluate literature review generation by LLMs have increasingly focused on assessing hallucinations in reference citations.
+For instance, Chelli et al.
+(2024) analyzed hallucination rates in 11 systematic reviews on shoulder rotator cuff pathology generated by ChatGPT, GPT-4, and Bard, finding Bard exhibited significantly higher hallucination rates.
 Similarly, Agrawal et al.
-(2024) chose 200 computer science topics to evaluate LLMs’ hallucination rates.
-They first asked LLMs to generate 5 reference titles, then used the BING Search API to automatically label the references as true or not.
-The LLMs were then examined to determine if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
+(2024) evaluated hallucinations across 200 computer science topics by generating reference titles with LLMs and verifying their existence via the BING Search API, further probing whether LLMs could detect hallucinated references through direct and indirect queries.
 Athaluri et al.
-(2023) assessed the hallucinations of LLMs in research proposal writing.
-They used ChatGPT to generate 50 research proposals based on 50 given topics.
-ChatGPT generated 178 references in these proposals, and then they employed five researchers to manually check these references and DOIs on Scopus, Google, and PubMed search engines.
-They found that among the references, 109 had valid DOIs.
-Moreover, Aljamaan et al.
-(2024) provided 5 medical topics and used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication date, link to the paper, and relevance of the reference to the keyword prompt.
-They then proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date.
-While these studies are very useful attempts at evaluating citation hallucinations in literature reviews, they are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance.
-Moreover, these evaluations often involve manual human intervention.
+(2023) examined hallucinations in 50 ChatGPT-generated research proposals, manually validating references and DOIs using Scopus, Google, and PubMed, reporting 109 valid DOIs among 178 references.
+Additionally, Aljamaan et al.
+(2024) introduced the Reference Hallucination Score (RHS) by generating references for five medical topics across multiple LLMs, assigning weighted hallucination scores to citation components such as title and publication date.
+While these studies provide valuable insights, they are generally limited to specific domains and rely heavily on manual evaluation, lacking a comprehensive, scalable assessment framework for LLM reference hallucinations.
 3 Methodology
-In this section, we propose a framework for evaluating LLMs’ literature survey writing ability.
+In this section, we propose a framework for evaluating LLMs’ literature review writing ability.
 The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
+Figure 1: Illustration of the evaluation framework.
 3.1 Dataset Construction
@@
 Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles.
-Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
+Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r_{1},...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
 3.2 Task Design
-Since literature review writing primarily involves the collection and synthesis of relevant research, we design three tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
-  • Task 1: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic.
-Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page.
+Since literature review writing primarily involves the collection and synthesis of relevant research, we design three independent tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
+  • Reference Generation: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic.
+Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page, $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$.
 In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
-  • Task 2: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract $a_{i}^{g}$according to the research topic provided in the title and the keywords.
-The length of the generated abstract is to be the same as the length of the original abstract of the article.
-In this task, we evaluate whether LLMs can write an abstract based on the given topic.
-  • Task 3: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract.
-The literature survey should be about 1000 words long.
-LLMs also need to back up claims by citing previous studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature survey).
-In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
+Note that these references are not reused in later tasks.
+  • Abstract Writing: Given an article title $t_{i}$ and its associated keywords $w_{i}$, the LLMs are prompted to generate an abstract $a_{i}^{g}$ that aligns with the research topic.
+The length of the generated abstract is constrained to match that of the original.
+This task serves as a proxy for literature review planning, as abstracts often outline the key components of a study—such as its objectives, methods, and covered subtopics—which are also critical in structuring comprehensive literature reviews.
+By evaluating the model’s ability to generate coherent and topic-relevant abstracts, we assess its potential to assist researchers in the early planning stages of literature review writing.
+  • Review Composition: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract.
+To facilitate evaluation and accommodate computational budget constraints, the length of each literature review is limited to approximately 1000 words.
+LLMs also need to back up claims by citing relevent studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature review).
+These citations are newly generated to support the content of the review.
+In this task, we evaluate whether LLMs can write a high-quality literature review and cite truth studies.
 Three task prompts are shown in Appendix Table LABEL:at1.
@@
 Based on the type of generated text, we divide the evaluation of the model’s results into two parts: first, the hallucination rate of the references generated by LLMs, and second, a comparison of the generated context with human-written results, including two dimensions: factual consistency and semantic coverage.
-Figure 1: Illustration of the experiment process.
 Reference hallucination evaluation metrics.
 Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references.
-In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM.
-A higher accuracy metric indicates a lower hallucination rate.
-For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task 1 and Task 3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the generated article title, authors, journal, year, volume, first page, and last page.
-Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
-Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$.
+In this section, we introduce the calculation process of the reference precision $Precison$, reference recall $Recall$, $F1$, and title search rate $S_{t}$ for each LLM.
+A higher precision metric indicates a lower hallucination rate.
+A higher recall indicates that the LLM-generated references cover more of the ground-truth citations used by human authors, reflecting a better ability to identify key prior work relevant to the topic.
+For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in both Reference Generation and Review Composition tasks, each $r_{n}^{g}$ and includes 7 elements.
+Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\}_{d=0}^{6}$, $e_{d}$ = 1 or 0.
+Next, we describe how to obtain $\{e_{d}\}_{d=0}^{6}$.
 First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines.
-Then, we obtain two candidate sets $Z_{t}$ and $Z_{a}$ respectively.
-Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $Z_{t}$ and $Z_{a}$.
-For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$.
-Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $z_{n}^{g}$ of $r_{n}^{g}$.
-Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $z_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1).
-On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article.
-Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable.
-On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article.
-Therefore, when the other three elements in $r_{n}^{g}$ are matching, we also consider it as a reliable article.
-$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
-For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2.
-Then, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(p_{i})$ across articles, as shown in Eq 3.
-$Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
-$Acc.=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
-On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness.
-In  Agrawal et al.
-(2024)’s work, they determined by ground-truth labels assigned using the Bing search API.
-Encouraged by this work, we also calculate the title search scores of generated references for each LLM, formulated as follows.
-$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}$ (4)
-$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (5)
-where $s_{p_{i}}$ is the number of the generated title that can be matched in the external scholar base.
-We search the generated title from an external academic search engine to estimate whether the title can be matched.
+This results in two sets of candidate articles, $Z_{t}$ and $Z_{a}$ respectively, We then merge the two sets and remove duplicates to obtain the final candidate set $Z=\{z_{1},...,z_{j},...,z_{J}\}$.
+Subsequently, we compare the generated $r_{n}^{g}$ with the article $z_{j}$ from candidate sets $Z$.
+For example, if the title of a candidate article $z_{j}$ matches the title of $r_{n}^{g}$, then $e_{0}=1$.
+Finally, we find the best candidate article based on the sum of $\{e_{d}\}_{d=0}^{6}$, and the one with the largest sum is the best candidate article $z_{j}$ of $r_{n}^{g}$.
+Then, we compare the alignment degree between the generated reference $r_{n}^{g}$ and the best-matching candidate article $z_{j}$ to determine whether $r_{n}^{g}$ corresponds to a real article (as shown in Eq.
+1).
+We consider $r_{n}^{g}$ to be reliable under either of the following two conditions:
+Title-based matching: If the title $T$ is correct (i.e., $e_{0}=1$), and at least one other metadata element (e.g., author, journal, year, etc.) also matches, the reference is deemed reliable.
+To allow for minor variations, we consider the title to be correct if it achieves a match rate of at least 80% with the ground-truth title—a threshold determined through human evaluation.
+Metadata-based matching: If the title $T$ is incorrect (i.e., $e_{0}=0$), we still consider the reference reliable if at least three of the remaining metadata elements (author, journal, year, volume, first page, last page) match those of a real article.
+This allows us to identify true references even when the title is noisy or incomplete.
+$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vspace{-2mm}$ (1)
+For each paper $p_{i}$ in the dataset, we compute the reference precision of the LLM-generated references, denoted as $\textit{Precision}{(p_{i})}$, as defined in Eq.
+2.
+We then obtain the overall Precision score for each LLM by averaging $\textit{Precision}{(p_{i})}$ across all papers in the dataset, as shown in Eq.
+3.
+$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vspace{-2mm}$ (2)
+$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vspace{-1mm}$ (3)
+Precision is measured by comparing the LLM-generated references with external academic databases.
+We also evaluate Recall by comparing the references generated by the LLM with those cited in the human-written original articles.
+The key difference between precision and recall in our setting lies in the candidate set $Z$: for precision, $Z$ is constructed from external academic search results, whereas for recall, $Z$ consists of the references actually cited in the human-written articles.
+Based on precision and recall, we further compute the F1 score, which serves as a harmonic mean to reflect the overall accuracy of the reference generation.
+Additionally, the title is intuitively the most critical element in determining the faithfulness of a generated reference.
+In the work of Agrawal et al.
+(2024), ground-truth labels were assigned based on results returned by the Bing Search API.
+Inspired by their approach, we also calculate the title search score for each LLM to estimate how many generated titles correspond to real publications.
+$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vspace{-2mm}$ (4)
+$s_{p_{i}}^{(n)}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}\vspace{-1mm}$ (5)
+Here, $s_{p_{i}}^{(n)}$ indicates whether the generated title in reference $r_{n}^{g}$ for paper $p_{i}$ can be found using an external academic search engine.
+This metric helps estimate the proportion of references with verifiable titles among the total generated references.
 Context evaluation metrics.
-In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects.
+In our study, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects.
 The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al.
@@
 where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 7.
-$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}$ (7)
+$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vspace{-2mm}$ (7)
 Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context.
-Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation.
-For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
+Specifically, for the Abstract Writing task, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation.
+For the Review Composition task, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
 KPR, first proposed by  Qi et al.
@@
 8.
-$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (8)
+$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vspace{-2mm}$ (8)
 where $c_{i}^{g}$ denotes the literature review generated by LLMs.
@@
 4.1 Experimental Settings
-Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
-LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
-In task 1 and task 3, we set $N$ as 10, each model generates 10 references.
+Dataset. We collect 1,105 literature review articles published in 2023 from the Annual Reviews website. The distribution of articles across journals is shown in Appendix B, Figure 4.
+LLMs Selection.
+We evaluate five LLMs: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
+All model outputs were generated via their official APIs with temperature set to 0 for consistency.
+In Reference Generation and Review Composition tasks, we set $N$ as 10, each model generates 10 references.
 For the generated reference evaluation, we use Semantic Scholar as the external database.
@@
 (2024).
-So, for Task 2, we employ TRUE Honovich et al.
+So, for the Abstract Writing task, we employ TRUE Honovich et al.
 (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; to compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings.
-For task 3, we employ GPT-4o as the NLI model in Eq 8, and set $q$ as 10.
+For the Review Composition task, we employ GPT-4o as the NLI model in Eq 8, and set $q$ as 10.
 4.2 Main Results
 We present the results for the three tasks in Table 12, and 3. The performance of different models on each task is analyzed as follows.
-Results for Task 1.
-By analyzing the results in Table 1, we observe that Claude-3.5-Sonnet achieves the highest accuracy and title search rate, whereas Llama-3.2-3B exhibits the lowest performance on both metrics for Task 1.
-Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models.
-This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs.
-Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2.
-The graph shows that the majority of the correctly cited references for almost all LLMs involved in the task are concentrated between 2010 and 2020.
-Results for Task 2.
-We use various metrics to evaluate model performance.
-As shown in Table 2, Claude-3.5-Sonnet consistently achieves the best overall performance across nearly all metrics.
-The average semantic similarity between the abstracts generated by Claude-3.5-Sonnet and human-written abstracts is 81.17%.
-Furthermore, Claude-3.5-Sonnet and DeepSeek-V3 exhibit the highest factual consistency, achieving a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.84%, respectively.
-Additionally, we observe that Llama-3.2-3B achieves the highest score on the ROUGE-L metric but does not exhibit a clear advantage on other metrics.
-This suggests that, given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
-Results for Task 3.
-As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3.
-Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text (Gao et al., 2023).
-In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly.
-This suggests that the generated references and the literature review text impose constraints on each other.
-On the other hand, we find that Claude-3.5-Sonnet achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5-Sonnet recalls the most claims from the human-written literature review.
-While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
-  Models Task 1
-$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
-Qwen-2.5-72B 21.80 12.25 17.58
-Llama-3.2-3B 16.62 3.45 6.95
-DeepSeek-V3 56.04 46.33 50.66
-GPT-4o 32.07 21.65 24.65
-Claude-3.5-Sonnet 64.82 51.59 55.77
-Table 1: The experimental results of the five LLMs in Task 1.
-“$S_{t}$” refers to the title search rate as defined in Eq 4, while “Acc.” represents the accuracy defined in Eq 3.
+Results for Reference Generation.
+As shown in Table 1, Claude-3.5-Sonnet achieves the highest $F1$ score and $S_{t}$, while Llama-3.2-3B performs the worst on both metrics.
+Notably, Claude-3.5-Sonnet also achieves the highest precision and recall, indicating that it not only generates more correct references overall but also shares the highest overlap with those cited in the human-written article.
+When evaluating the author dimension of LLM-generated references, we consider the reference to match in this dimension if the first author is correctly matched.
+Applying this criterion results in a 1–3% increase in F1 scores across all models.
+This suggests that generating complete and accurate author lists remains a major challenge for LLMs.
+We further conduct a year-wise analysis of the correctly generated references, as illustrated in Figure 2.
+The results reveal that the majority of accurate citations produced by the models are concentrated in the period between 2010 and 2020, a trend consistent across nearly all LLMs evaluated in this task.
+Results for Abstract Writing.
+As shown in Table 2, Claude-3.5-Sonnet achieves the best overall performance across most evaluation metrics.
+It generates abstracts with the highest average semantic similarity to human-written ones (81.17%) and shows strong factual consistency, achieving a TRUE score of 78.10%.
+DeepSeek-V3 also performs well in factual consistency, with the highest GPT-4o-based assessment score (96.84%).
+In contrast, Llama-3.2-3B obtains the highest ROUGE-L score but does not show clear advantages on other metrics.
+These results highlight the importance of using multiple evaluation metrics to comprehensively assess the diverse outputs of LLMs.
+Results for Review Composition.
+As shown in Table 3, compared to the Reference Generation task, all LLMs demonstrate a significant increase in Precision when generating references within the Review Composition task.
+Prior research indicates that grounding generated text with real external citations can effectively reduce hallucination rates (Gao et al., 2023).
+Consistently, our experiments reveal that when LLMs generate references alongside the literature review, the accuracy of these references improves markedly.
+This suggests a mutual constraint between the generated references and the review text, leading to enhanced overall reliability.
+On the other hand, Claude-3.5-Sonnet achieves the highest performance on the KPR metric, indicating that its generated literature reviews recall the greatest number of claims from the human-written versions.
+Meanwhile, the literature reviews produced by DeepSeek-V3 excel on the ROUGE metrics, demonstrating stronger overlap with reference texts in terms of lexical similarity.
+  Models Reference Generation
+$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ P R F1 P R F1(first author)
+Qwen-2.5-72B 21.80 12.25 12.60 12.42 17.58 13.27 15.12
+Llama-3.2-3B 16.62 3.45 8.48 4.90 6.95 8.67 7.72
+DeepSeek-V3 56.04 46.33 19.72 27.66 50.66 20.50 29.19
+GPT-4o 32.07 21.65 18.76 20.10 24.65 19.50 21.77
+Claude-3.5-Sonnet 64.82 51.59 24.34 33.08 55.77 25.21 34.72
+Table 1: The experimental results of the five LLMs in Reference Generation.
+“$S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall.
 “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
-  Models Task 2
+  Models Abstract Writing
 […]
@@
 Claude-3.5-Sonnet 81.17 78.90 96.77 41.13 8.99 20.00
-Table 2: Compare the performance of four LLMs on Task 2.
-  Models Task 3
+Table 2: Compare the performance of four LLMs on Abstract Writing.
+  Models Review Composition
 References Literature Review
 […]
-Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
-Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
-DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
-GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
-Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
-Table 3: The experimental results of the four LLMs in Task 3.
-“$S_{t}$” refers to the title retrieval rate as defined in Eq 4.“Acc.” represents the accuracy defined in Eq 4.
+Qwen-2.5-72B 40.02 28.91 17.36 21.69 33.64 18.31 23.713 38.82 29.95 9.01 15.14
+Llama-3.2-3B 21.78 4.86 8.28 6.12 7.28 8.44 7.82 29.07 28.07 7.77 15.46
+DeepSeek-V3 62.29 52.81 26.79 35.55 55.38 27.30 36.57 56.02 35.65 10.40 17.46
+GPT-4o 60.05 50.62 27.88 35.96 54.16 28.86 37.65 59.18 30.78 9.72 15.54
+Claude-3.5-Sonnet 66.43 59.06 31.90 41.42 63.06 33.25 43.54 62.32 28.59 8.90 14.41
+Table 3: The experimental results of the four LLMs in Review Composition.
+“$S_{t}$” refers to the title retrieval rate as defined in Eq 4.
+While “P” represents the Precision, “R” denotes the Recall.
 “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
-  Models Task1 Task3
-Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
-Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
-Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
-DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
-GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
-Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
-Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3.
-“Title Acc.” refers to the title overlap between LLM-generated and human-written references, while “overlap” represents the article overlap between LLM-generated and human-written references.
-“first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
+  Discipline Citation Count Precision
+DeepSeek Claude DeepSeek Claude
+Biology 763 678 55.55 58.00
+Mathematics 2288 1984 60.00 62.22
+Physics 894 652 47.62 56.19
+Chemistry 1334 1079 43.14 43.80
+Social Science 1321 1151 46.80 56.70
+Technology 904 748 44.88 49.01
+Table 4: Average citation counts and reference precision across disciplines.
 Figure 2: Distribution of LLM-generated true references over years.
 4.3 Analyze LLM-Generated References from Different Dimensions
-In both Task 1 and Task 3, we ask LLMs to generate references.
+In both Reference Generation and Review Composition tasks, we ask LLMs to generate references.
 The overall performance was discussed in the previous section.
 In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2.
-As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in Task 1.
-Additionally, the accuracy of reference generation in Task 1 for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
+As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in the Reference Generation task.
+Additionally, the accuracy of reference generation in the Reference Generation task for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension.
 Accuracy for journal name, page, and author is also relatively high.
@@
 However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
-Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension.
+Next, we examine the accuracy of LLM-generated references across various dimensions in Review Composition, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension.
 The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension.
@@
 We then present bar charts in Figures  LABEL:f4, which illustrate the performance of different models across various tasks and disciplines.
-First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline.
-Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences.
-The results are as follows: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
-These results indicate significant differences in performance across disciplines for each model except Llama-3.2-3B.
-Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2.
-We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology.
-Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines.
-The results are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
-These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
-Additionally, we examine the references accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c.
-It is evident that the accuracy of Claude-3.5-Sonnet, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines.
-Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology.
-The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001), indicating significant differences in each models’ performance across the various disciplines.
-Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d.
-The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests.
-Specifically, the p-values are as follows: Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
-4.5 Comparison of LLM-Cited and Human-cited References
-In this section, we also compare the LLM-generated references with those from the human-written original articles.
-Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article.
-As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5-Sonnet are also cited in the human-written article, indicating a relatively high overlap rate.
-Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3.
-Notably, GPT-4o models show a 10% increase in the overlap rate.
-We provide a more detailed comparison of the references across various dimensions in Appendix C.
+First, we observe that in the Reference Generation task, as shown in Figure LABEL:f4a, almost all models exhibit the highest precision in the Mathematics discipline and the lowest precision in the Chemistry discipline.
+To validate these differences, we conduct one-way ANOVA tests for each LLM across five disciplines.
+Significant differences are found for all models except Llama-3.2-3B.
+Detailed ANOVA results are reported in Appendix G.
+Secondly, as shown in Figure LABEL:f4b, the NLI scores evaluated by TRUE in the Abstract Writing task indicate that all models perform the worst in Social Science.
+GPT-4o performs best in Technology, while Claude 3.5-Sonnet achieves the highest performance in Biology.
+One-way ANOVA tests reveal significant differences across disciplines for all models.
+See Appendix G for detailed results.
+Thirdly, we examine the references precision of each model across five disciplines in the Review Composition task, as illustrated in Figure LABEL:f4c.
+It is evident that the precision of Claude 3.5 Sonnet, DeepSeek-V3, and GPT-4o is significantly higher than that of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines.
+Furthermore, Claude 3.5 Sonnet, DeepSeek-V3, and Qwen-2.5-72B exhibit the highest precision in Mathematics, while GPT-4o performs best in Social Science.
+ANOVA tests confirm significant differences across disciplines for all models (see Appendix G).
+Finally, we observe the KPR metric across different disciplines in Review Composition, as shown in Figure LABEL:f4d.
+The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests (see Appendix G).
+Citation Frequency and Precision Across Disciplines.
+Additionally, we report statistics on the citation frequency of correctly generated references by LLMs in the Reference Generation task, as shown in Table 4, using Claude-3.5 and DeepSeek-V3 as examples.
+The data indicates that the references generated by the LLMs are highly cited, which might be due to their frequent presence in online sources, making them more likely to appear in the LLMs training datasets.
+As a result, LLMs tend to generate more accurate metadata (e.g., author, year) for these well-known references.
+Furthermore, when analyzing different disciplines, we observe that the Mathematics discipline has the highest precision, and the relevant references generated for Mathematics also have the highest citation count.
+We compute the correlation between citation precision and average citation counts, finding that the correlation coefficient for Claude-3.5 is 0.4, and for DeepSeek-V3 it is 0.51, indicating a positive relationship between the two.
 4.5 Human Evaluation
@@
 Limitations
-In this paper, we evaluate the ability of LLMs to write literature reviews.
-However, instead of assessing from conventional perspectives, such as the fluency of the generated context or topic coverage, we compare the model-generated results with human-written ones.
-Additionally, when processing the LLM-generated results, we often need to handle abbreviations of author names and journal titles.
-While we have made efforts to address these details thoroughly, some discrepancies may still exist.
+In this paper, we evaluate the ability of LLMs to write literature reviews. However, several limitations remain:
+First, instead of evaluating the generated reviews from conventional perspectives such as fluency or topic coverage, we primarily compare LLM-generated results with human-written ones.
+As such, our current evaluation metrics may not be comprehensive.
+In the future, we plan to incorporate additional aspects of review quality to improve the completeness of our evaluation.
+These may include the coverage of cited works (i.e., whether the review offers a comprehensive overview of the relevant field) and the coherence of the overall structure (i.e., whether the review is organized in a way that facilitates information-seeking).
+Second, there is a possibility that our test data overlaps with the training data of the LLMs.
+When we initiated this study in August 2024, the dataset from the Annual Reviews website had not yet been updated to include 2024 articles, so we relied on the complete 2023 dataset.
+To mitigate potential data leakage, we plan to deploy a leaderboard on Hugging Face to continuously evaluate the performance of various LLMs in literature review writing, with real-time updates to the test dataset.
+However, due to the rapid iteration of LLMs, data leakage cannot be completely ruled out.
+That said, our experimental results—particularly those related to reference generation—show that all models still perform poorly.
+If data contamination were present, the actual scores would likely be lower than those reported.
+This reinforces, rather than undermines, our conclusion that significant challenges remain in using LLMs for literature review generation.
+Additionally, when processing LLM-generated outputs, we frequently encountered abbreviated author names and journal titles.
+Although we have taken care to address these issues thoroughly (see Appendix E), minor discrepancies may still exist.
+Finally, to verify the precision of LLM-generated references, we primarily used Semantic Scholar as our auxiliary tool.
+Although we also experimented with Google Scholar, its lack of an accessible API led us to rely on the freely available Semantic Scholar API for consistency and ease of access.
+However, this may have resulted in incomplete reference retrieval.
+Ethics Statement
+The human evaluations conducted in this study were carried out by members of the research team.
+No personal or sensitive information was collected, and all participants were fully informed of the purpose of the evaluation.
+Therefore, the study does not raise any ethical concerns.
 Appendix A Prompts for Tasks
@@
 We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions.
-As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower.
-For Task 3, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1.
-This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
+As shown in Figure LABEL:f3, for Reference Generation, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower.
+For Review Composition, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Reference Generation.
+This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Review Composition leads to the generation of more accurate author information.
 Appendix D Discussion
@@
 However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
-When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references.
+When comparing the references generated by the models in Reference Generation and Review Composition, we find that in Review Composition, nearly all models generate more accurate references.
 This suggests that LLMs cite references during the writing process, which improves the authenticity of the references.
@@
 In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.
+Appendix E Data Processing Strategy
+Author name and journal title variations often pose challenges when aligning LLM-generated references with articles from Semantic Scholar.
+To address this, we adopt the following normalization strategies:
+Author names.
+When comparing author names between LLM-generated references and candidate articles, if an exact match is not found (e.g., “John Smith”), we consider common variants such as “Smith, John”, “Smith, J.”, or “J.
+Smith” to account for different citation formats.
+Journal titles.
+For journal names like Journal of Chemical Physics, we incorporate standard abbreviation forms (e.g., J.
+Chem.
+Phys.) based on widely used abbreviation conventions.
+Nonetheless, certain non-standard or ambiguous cases may still be unmatched.
+Appendix F LLM Output Example
+Here is one example of the output generated by Claude-3.5-Sonnet for the literature review composition task.
+{"Literature Review": "CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules.
+This review explores the technical developments and application prospects of CRISPR-based biosensing strategies, highlighting their potential to revolutionize diagnostics and molecular detection.The core biochemical properties of CRISPR-Cas systems provide the foundation for their application in biosensing.
+These properties include diverse reaction temperatures, programmability in design, high reaction efficiency, and recognition specificity.
+Chen et al.
+(2018) demonstrated the versatility of CRISPR-Cas9 in detecting DNA targets across a wide temperature range, from 20°C to 60°C, enabling its use in various environmental conditions.
+The programmability of CRISPR systems allows for the easy design of guide RNAs (gRNAs) to target specific sequences, as shown by Gootenberg et al.
+(2017) in their development of the SHERLOCK (Specific High-sensitivity Enzymatic Reporter unLOCKing) platform for nucleic acid detection.The high reaction efficiency of CRISPR-Cas systems contributes to their sensitivity in biosensing applications.
+Li et al.
+(2019) reported a CRISPR-Cas12a-based assay capable of detecting attomolar concentrations of DNA targets, demonstrating the potential for ultra-sensitive detection.
+The recognition specificity of CRISPR systems allows for precise discrimination between closely related sequences, as evidenced by the work of Myhrvold et al.
+(2018), who developed a CRISPR-based diagnostic platform capable of distinguishing between Zika and Dengue virus strains with single-base resolution.Recent efforts have focused on improving these core properties to enhance the performance of CRISPR-based biosensors.
+For example, Harrington et al.
+(2018) engineered Cas12a variants with improved activity and specificity, expanding the range of targetable sequences and reducing off-target effects.
+Similarly, Gootenberg et al.
+(2018) developed an enhanced version of the SHERLOCK system, incorporating lateral flow readouts and multiplexing capabilities, further improving its utility in point-of-care diagnostics.One of the key areas of technical development in CRISPR-based biosensing is improving sensitivity and quantification capabilities.
+Wang et al.
+(2020) introduced a CRISPR-Cas12a-based digital PCR method that achieved attomolar sensitivity and precise quantification of target nucleic acids.
+This approach combines the specificity of CRISPR recognition with the quantitative power of digital PCR, opening new avenues for highly sensitive and accurate molecular diagnostics.Multiplex detection is another important aspect of CRISPR biosensing development.
+Ackerman et al.
+(2020) demonstrated a CRISPR-Cas13-based multiplexed assay capable of simultaneously detecting and distinguishing multiple SARS-CoV-2 variants, showcasing the potential of CRISPR technology in rapid and accurate pathogen identification.
+The ability to detect multiple targets in a single reaction not only saves time and resources but also provides more comprehensive diagnostic information.Efforts to develop convenient one-pot assays have led to simplified and more user-friendly CRISPR-based detection methods.
+Joung et al.
+(2020) reported a one-pot, CRISPR-Cas12-based assay for SARS-CoV-2 detection that could be performed at a single temperature, eliminating the need for complex thermal cycling equipment.
+This advancement brings CRISPR-based diagnostics closer to point-of-care applications, where simplicity and ease of use are crucial.The creation of advanced sensors has further expanded the capabilities of CRISPR-based biosensing.
+Bruch et al.
+(2019) developed a CRISPR-Cas13a-based electrochemical sensor for nucleic acid detection, demonstrating the potential for integrating CRISPR technology with electronic readout systems.
+This approach offers the possibility of developing portable, electronic CRISPR-based diagnostic devices for rapid and sensitive detection in resource-limited settings.While CRISPR-based biosensing has shown great promise, there are still obstacles to its widespread commercial application.
+These include challenges in sample preparation, the need for standardization across different CRISPR systems and assay formats, and regulatory considerations for diagnostic use.
+Addressing these issues will be crucial for the successful translation of CRISPR biosensing technology from the laboratory to real-world applications.Looking ahead, the future of CRISPR-based biosensing appears bright, with numerous opportunities for further development and application.
+Potential directions include the integration of CRISPR biosensors with microfluidic devices for automated sample processing and analysis, the development of CRISPR-based wearable sensors for continuous health monitoring, and the application of machine learning algorithms to improve the design and performance of CRISPR biosensing systems.In conclusion, CRISPR-based biosensing strategies represent a rapidly evolving field with immense potential to transform molecular detection and diagnostics.
+As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.", "References": [{"title": "CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity", "authors": "Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "436", "last page": "439"},{"title": "Nucleic acid detection with CRISPR-Cas13a/C2c2", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.", "journal": "Science", "year": "2017", "volumes": "356", "first page": "438", "last page": "442", "DOI": "10.1126/science.aam9321"}, {"title": "CRISPR-Cas12a-assisted nucleic acid detection", "authors": "Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin", "journal": "Cell Discovery", "year": "2019", "volumes": "5", "first page": "1", "last page": "4"}, {"title": "Field-deployable viral diagnostics using CRISPR-Cas13", "authors": "Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "444", "last page": "448"}, {"title": "Enhanced proofreading governs CRISPR–Cas9 targeting accuracy", "authors": "Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.", "journal": "Nature", "year": "2018", "volumes": "563", "first page": "621", "last page": "625"}, {"title": "Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng", "journal": "Science", "year": "2018", "volumes": "360", "first page": "439", "last page": "444"}, {"title": "Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay", "authors": "Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua", "journal": "Nature Communications", "year": "2020", "volumes": "11", "first page": "4711", "last page": "4711"}, {"title": "Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing", "authors": "Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P.
+J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem", "journal": "New England Journal of Medicine", "year": "2020", "volumes": "383", "first page": "1492", "last page": "1494"}, {"title": "CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2", "authors": "Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.", "journal": "Angewandte Chemie International Edition", "year": "2019", "volumes": "58", "first page": "17571", "last page": "17575"}, {"title": "Scalable and robust SARS-CoV-2 testing in an academic center", "authors": "Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared", "journal": "Nature Biotechnology", "year": "2020", "volumes": "38", "first page": "927", "last page": "931"}]}
+Appendix G ANOVA Test Results
+We report the p-values from one-way ANOVA tests across disciplines for each model and task:
+  • Reference Generation task: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
+  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
+  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001).
+  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).

### v5

@@
 This allows us to identify true references even when the title is noisy or incomplete.
-$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vspace{-2mm}$ (1)
+$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vskip-5.69054pt$ (1)
 For each paper $p_{i}$ in the dataset, we compute the reference precision of the LLM-generated references, denoted as $\textit{Precision}{(p_{i})}$, as defined in Eq.
@@
 3.
-$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vspace{-2mm}$ (2)
-$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vspace{-1mm}$ (3)
+$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vskip-5.69054pt$ (2)
+$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vskip-2.84526pt$ (3)
 Precision is measured by comparing the LLM-generated references with external academic databases.
@@
 Inspired by their approach, we also calculate the title search score for each LLM to estimate how many generated titles correspond to real publications.
-$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vspace{-2mm}$ (4)
-$s_{p_{i}}^{(n)}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}\vspace{-1mm}$ (5)
+$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vskip-5.69054pt$ (4)
+[…]
 Here, $s_{p_{i}}^{(n)}$ indicates whether the generated title in reference $r_{n}^{g}$ for paper $p_{i}$ can be found using an external academic search engine.
@@
 where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 7.
-$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vspace{-2mm}$ (7)
+$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vskip-5.69054pt$ (7)
 Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context.
@@
 8.
-$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vspace{-2mm}$ (8)
+$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vskip-5.69054pt$ (8)
 where $c_{i}^{g}$ denotes the literature review generated by LLMs.
@@
 Table 1: The experimental results of the five LLMs in Reference Generation.
-“$S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall.
+$S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall.
 “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@
 Table 3: The experimental results of the four LLMs in Review Composition.
-“$S_{t}$” refers to the title retrieval rate as defined in Eq 4.
+$S_{t}$” refers to the title retrieval rate as defined in Eq 4.
 While “P” represents the Precision, “R” denotes the Recall.
@@
 Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature.
-I would like you to find the 10 studies that are most relevant to the research topic provided in the "Title" and the "Keywords" below.
+I would like you to find the 10 studies that are most relevant to the research topic provided in the ”Title” and the ”Keywords” below.
 Please cite the studies according to the following JSON format.
 There is no need to provide any explanation before or after the JSON output.
-Ensure that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
-{ "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
+Ensure that the ”authors” field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as ”et al.”.
+{ ”References”: [ { ”title”: ””, ”authors”: ””, ”journal”: ””, ”year”: ””, ”volumes”: ””, ”first page”: ””, ”last page”: ””, } ] }
 Title: title
@@
 Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature.
-I would like you to write an abstract according to the research topic provided in the "Title" and the "Keywords" below.
+I would like you to write an abstract according to the research topic provided in the ”Title” and the ”Keywords” below.
 Please write the abstract for about xx words, according to the JSON format as follows.
 There is no need to provide any explanation before or after the JSON output.
-{"Abstract": ""}
+{”Abstract”: ””}
 Title: title
@@
 Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature.
-I would like you to write a literature review according to the research topic provided in the "Title", “Abstract” and "Keywords" below.
+I would like you to write a literature review according to the research topic provided in the ”Title”, “Abstract” and ”Keywords” below.
 The literature review should be about 1000 words long.
 I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review).
-The output should be in JSON format as follows: { "Literature Review": "xxx", "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
-The "Literature Review" field should be about 1000 words.
-The "References" field is a list of 10 references, and ensures that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
+The output should be in JSON format as follows: { ”Literature Review”: ”xxx”, ”References”: [ { ”title”: ””, ”authors”: ””, ”journal”: ””, ”year”: ””, ”volumes”: ””, ”first page”: ””, ”last page”: ””, } ] }
+The ”Literature Review” field should be about 1000 words.
+The ”References” field is a list of 10 references, and ensures that the ”authors” field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as ”et al.”.
 Title: title
@@
 Here is one example of the output generated by Claude-3.5-Sonnet for the literature review composition task.
-{"Literature Review": "CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules.
+{”Literature Review”: ”CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules.
 This review explores the technical developments and application prospects of CRISPR-based biosensing strategies, highlighting their potential to revolutionize diagnostics and molecular detection.The core biochemical properties of CRISPR-Cas systems provide the foundation for their application in biosensing.
@@
 Potential directions include the integration of CRISPR biosensors with microfluidic devices for automated sample processing and analysis, the development of CRISPR-based wearable sensors for continuous health monitoring, and the application of machine learning algorithms to improve the design and performance of CRISPR biosensing systems.In conclusion, CRISPR-based biosensing strategies represent a rapidly evolving field with immense potential to transform molecular detection and diagnostics.
-As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.", "References": [{"title": "CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity", "authors": "Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "436", "last page": "439"},{"title": "Nucleic acid detection with CRISPR-Cas13a/C2c2", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.", "journal": "Science", "year": "2017", "volumes": "356", "first page": "438", "last page": "442", "DOI": "10.1126/science.aam9321"}, {"title": "CRISPR-Cas12a-assisted nucleic acid detection", "authors": "Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin", "journal": "Cell Discovery", "year": "2019", "volumes": "5", "first page": "1", "last page": "4"}, {"title": "Field-deployable viral diagnostics using CRISPR-Cas13", "authors": "Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "444", "last page": "448"}, {"title": "Enhanced proofreading governs CRISPR–Cas9 targeting accuracy", "authors": "Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.", "journal": "Nature", "year": "2018", "volumes": "563", "first page": "621", "last page": "625"}, {"title": "Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng", "journal": "Science", "year": "2018", "volumes": "360", "first page": "439", "last page": "444"}, {"title": "Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay", "authors": "Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua", "journal": "Nature Communications", "year": "2020", "volumes": "11", "first page": "4711", "last page": "4711"}, {"title": "Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing", "authors": "Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P.
-J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem", "journal": "New England Journal of Medicine", "year": "2020", "volumes": "383", "first page": "1492", "last page": "1494"}, {"title": "CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2", "authors": "Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.", "journal": "Angewandte Chemie International Edition", "year": "2019", "volumes": "58", "first page": "17571", "last page": "17575"}, {"title": "Scalable and robust SARS-CoV-2 testing in an academic center", "authors": "Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared", "journal": "Nature Biotechnology", "year": "2020", "volumes": "38", "first page": "927", "last page": "931"}]}
+As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.”, ”References”: [{”title”: ”CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity”, ”authors”: ”Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”436”, ”last page”: ”439”},{”title”: ”Nucleic acid detection with CRISPR-Cas13a/C2c2”, ”authors”: ”Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.”, ”journal”: ”Science”, ”year”: ”2017”, ”volumes”: ”356”, ”first page”: ”438”, ”last page”: ”442”, ”DOI”: ”10.1126/science.aam9321”}, {”title”: ”CRISPR-Cas12a-assisted nucleic acid detection”, ”authors”: ”Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin”, ”journal”: ”Cell Discovery”, ”year”: ”2019”, ”volumes”: ”5”, ”first page”: ”1”, ”last page”: ”4”}, {”title”: ”Field-deployable viral diagnostics using CRISPR-Cas13”, ”authors”: ”Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”444”, ”last page”: ”448”}, {”title”: ”Enhanced proofreading governs CRISPR–Cas9 targeting accuracy”, ”authors”: ”Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.”, ”journal”: ”Nature”, ”year”: ”2018”, ”volumes”: ”563”, ”first page”: ”621”, ”last page”: ”625”}, {”title”: ”Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6”, ”authors”: ”Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”439”, ”last page”: ”444”}, {”title”: ”Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay”, ”authors”: ”Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua”, ”journal”: ”Nature Communications”, ”year”: ”2020”, ”volumes”: ”11”, ”first page”: ”4711”, ”last page”: ”4711”}, {”title”: ”Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing”, ”authors”: ”Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P.
+J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem”, ”journal”: ”New England Journal of Medicine”, ”year”: ”2020”, ”volumes”: ”383”, ”first page”: ”1492”, ”last page”: ”1494”}, {”title”: ”CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2”, ”authors”: ”Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.”, ”journal”: ”Angewandte Chemie International Edition”, ”year”: ”2019”, ”volumes”: ”58”, ”first page”: ”17571”, ”last page”: ”17575”}, {”title”: ”Scalable and robust SARS-CoV-2 testing in an academic center”, ”authors”: ”Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared”, ”journal”: ”Nature Biotechnology”, ”year”: ”2020”, ”volumes”: ”38”, ”first page”: ”927”, ”last page”: ”931”}]}
 Appendix G ANOVA Test Results
 We report the p-values from one-way ANOVA tests across disciplines for each model and task:
-  • Reference Generation task: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
-  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
-  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001).
-  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
+  • Reference Generation task: Claude-3.5-Sonnet (p¡.0001), DeepSeek-V3(p¡.0001), GPT-4o(p¡.0001), Qwen-2.5-72B (p¡.0001), Llama-3.2-3B (p=0.065).
+  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p¡.0001), DeepSeek-V3(p¡0.05), GPT-4o(p¡.01), Qwen2.5-72B (p¡.001), Llama-3.2-3B (p¡.0001).
+  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p¡.0001), DeepSeek-V3(p¡.0001), GPT-4o(p¡.0001), Qwen-2.5-72B(p¡.0001), Llama-3.2-3B(p¡.001).
+  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p¡0.001).